OpenStudy (anonymous):

A store sells a $39 shirt for $64.35. Find the percent of the markup.

hero (hero):

\[\frac{is}{of} = \frac{\%}{100} \]

hero (hero):

\[\frac{64.35}{39}=\frac{x}{100}\] Solving for x gives the percent markup

hero (hero):

165%

hero (hero):

Yeah, that makes more sense